Evaluate
-(-8/-4)^2-2-1^2=6

Answer by stanbon(75887)

-(-8/-4)^2-2-1^2=6
------
-(2)^2-2-1 = 6
-4-2-1 = 6
-7 = 6
----
That is what your equation says.
----
Cheers,
Stan H.
====